40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jan 25 16:30:16 2011, Tara, Frank, DailyProgress, Electronics Equipment, RCAV RFPD fixed RCAV-RFPD1.jpgRCAV-RFPD2.jpgnew_RCAV-RFPD3.jpg
    Reply  Fri Jan 28 13:35:35 2011, Tara, Frank, DailyProgress, Electronics Equipment, RCAV RFPD fixed 35MHzPD.pngpd21.png
       Reply  Sun Jan 30 22:58:35 2011, Tara, DailyProgress, Electronics Equipment, RCAV RFPD fixed beat_2011_01_30.png
Message ID: 457     Entry time: Sun Jan 30 22:58:35 2011     In reply to: 456
Author: Tara 
Type: DailyProgress 
Category: Electronics Equipment 
Subject: RCAV RFPD fixed 

I adjusted the temperature on ACAV and RCAV so that both cavities can be locked simultaneously.

I measured the beat noise and see improvement at low frequency after we fixed the RFPD.

 

After we reinstalling the RFPDs on the table, we need to wait for the temperature to settle before both cavities can be locked.

(The temperature readout jumps by ~0.05 degree if we work on the table. The cause is not known yet. we tried pulling the cables for T readout,

but this does nothing to the T readout.  This problem will be investigated soon.)

I check the beat noise and see some improvements at low frequency . It does not look a lot for now, but I think it might go lower

if I re align the beam to the cavities and the beam for beat measurement. It might drift away a bit during the break, and I haven't checked that yet.

 

The input power is ~ 1 mW for both cavities.

RCAV gain = 25

ACAV gain= 8.0 on the knob.

 

 

Quote:

The TF of 2 35.5MHz RFPD and 21.5 MHz RFPD are measured by the Jenne laser. RCAV's RFPD has the better response, while ACAV's RFPD peak is slightly shift to 37 MHz. But the peak is low Q so we lose only a few dB. PMC's RFPD is now tune to have a resonance peak at 21.5 MHz and a notch at 43 MHz as it should be.

 

 

Attachment 1: beat_2011_01_30.png  47 kB  | Hide | Hide all
beat_2011_01_30.png
ELOG V3.1.3-